An archer shoots an arrow at a 75 m distant target; the bull’s-eye of the target is at same height as the release height of the arrow. If d = 75 m and v = 37 m/s. At what angle must the arrow be released to hit the bull’s-eye if its initial speed is 37m/s?
1
Expert's answer
2021-07-19T15:42:39-0400

X- direction

Newton's second law

"S_x=ut+\frac{1}{2}at^2\\a=0
\\S_x=ut\\75=37(cos\theta)t"

Y-direction


"S_y=u_yt+\frac{1}{2}at^2"


"S_y=0,a=-g,u_y=usin\theta=37sin\theta"

"0=37(sin\theta)t+\frac{1}{2}(-9.8)t^2"

"0=(37sin\theta)+\frac{1}{2}(-9.8)t\\t=709.81sin\theta\\"

"75=37cos\theta \times(709.81\times sin\theta)\\75=24509.8sin\theta cos\theta\\"

"sin2\theta=2sin\theta cos\theta"

"75=13131.4851sin2\theta"

"\theta=16.5°"